OpenStudy (anonymous):

A line passes through the point (-4, -2) and has a slope of 3/2. What is the equation of the line? y= -3/2 x+4 y= 3/2 x+4 y= -3/2 x- 1/4 y= 1/2 x- 1/4

OpenStudy (anonymous):

Solve (y+2)/(x+4)=3/2 for y. #2

OpenStudy (anonymous):

Can you help me with another equation. What is the equation of the line that passes through (2, -7) and (-1, 2)? y = -3x - 1 y = 3x + 1 y = -1/3 x - 1 y = 1/3 x +1

OpenStudy (anonymous):

@robtobey